What is the probability that a seven-digit phone number contains the number 7?
Answer by Nate(3500)

possible combinations ~> 10^7
possible combinations of having any 7s ~> P(7,1) + P(7,2) + P(7,3) + P(7,4) + P(7,5) + P(7,6) + P(7,7) = 7 + 42 + 210 + 840 + 2520 + 5040 + 5040 = 13,699
Prob: 13,699 / 10^7 = 0.13699%